Vf = 10 ms-¹+2ms-²×5sec . I know it is from equations of motion but I have a question: Vi is 10 Ms minus 1 but when acceleration will multiply with time it will become 10 Ms minus 2 so how can Vi and acceleration be added. They have different powers.
1
Expert's answer
2018-02-26T09:54:07-0500
The unit of at is
[at]=[a][t]=m/s^2 s=m/s.
But this is the unit of velocity! Thus, they have the same powers.